# Object-Centric Latent Action Learning

### v2

@@ Appendix A Dataset collection
Distracting Control Suite (DCS): The datasets were collected via experts policies trained on DCS via PPO (for cheetah-run, walker-run and hopper-hop) and SAC (for humanoid-walk). The scores of the experts are presented on the Table 2. The final dataset of transitions for each tasks consists of 5k trajectories (1k transitions each). The observations in the dataset have height and width 64px.
Distracting MetaWorld (DMW): The datasets were collected via experts policies from MetaWorld. The scores of the experts are presented on the Table 2. The final dataset of transitions for each tasks consists of 20k successfull trajectories (we filtered out non-successfull trajectories during dataset collection). The observations in the dataset have height and width 128px.
@@ Appendix B Training details
All experiments were run in H100 GPU 80GB. The models are trained in bfloat16 precision. Training duration is shown in Table 4. We perform hyperparameter optimization for each model. The 3 best sets of parameters were evaluated and the best mean score across 3 seeds is reported in Figure 4.
Object-centric learning pretraining codebase was adopted from Videosaur Zadaianchuk et al. (2023). It utilizes DINOv2 (Oquab et al., 2023) pretrained encoder vit-base-patch14-dinov2.lvd142m from Timm (Wightman, 2019) models hub. The images in the dataset are upscaled for dino encoder up to 518px. The hyperparameters for object-centric pretraining can be found in Table 5
Latent action learning model for images and object-centric masks is formed from IDM and FDM models based on a ResNet-like CNN encoder and decoder. The hyperparameter optimization setups for latent action learning from images (used for lapo, lapo-masks) can be found in Table 7.
Latent action learning model for object-centric slots is formed from IDM and FDM based on 3-layer MLP blocks with residual connections and GeLU activations to effectively process vector representations. The hyperparameter optimizations setups for latent action learning from representations (used for lapo-slots) can be found in Table 7.
Appendix C Fixed Initialization for Slot Stability
[…]
@@ Appendix C Fixed Initialization for Slot Stability
Method Training Duration DCS Training Duration DMW
@@ Appendix D Evaluation scores on DCS-hard
We mark as DCS-hard, the experiments on Disctracting Control Suite with 3 types of distractions: dynamic backgroungs, color and camera position perturbations. The scale of color and camera variations is $0.1$. The aggregated mean scores for DCS-hard are present on Figure 1 and Table 1.
The corresponding results on DCS (only dynamic backgrounds) are presented on Figures 4 and 1 and Table 1.
@@ Appendix E Correlation between probe and normalized success rates
We examine the correlation between the linear action probe (described in Section 3) and the evaluation performance, using the same budget of labeled trajectories for both probe training and supervised fine-tuning of object-centric latent action learning (LAPO-slots), on the basketball task from DMW.
As shown in Figure 8, there is a strong correlation between the probe performance and the resulting success rate, suggesting that the probe can serve as a reliable method for slot selection.
Figure 8: Linear action probes and corresponding resulting normalized success rates of BC agents on same amounts of labeled trajectories for different number of slots (top-2 slots for $K\leq 5$, top-3 slots for $K\geq 8$). From left to right: (1) 32 labeled trajectories, (2) 64 labeled trajectories, (3) 128 labeled trajectories
@@ Appendix F Probe slots selection for DCS
Slot selection for tasks from Distracting Control Suite. The selected slots were used to obtain results for Figure 4 and Table 1.
@@ Appendix G Probe slots selection for DMW
Slot selection for tasks from Distracted MetaWord. The selected slots (ranking $I$) were used to obtain results for Figure 4 and Table 1.
@@ Appendix H Probe slots selection for different number of slots
Slot selection for varying number of slots for basketball task from Distracted MetaWord. The selected slots (ranking I) from Figures 18, 15, 19, 20, 21 and 22 were used to obtain the results for Figure 6.
@@
Figure 17: Slot selection for 2 slots. From left to right: (1) Linear action probes for VideoSaur slots pretrained basketball task for varying budget of trajectories. (2) Corresponding slot masks for basketball task after VideoSaur pretraining for number of slots $K=2$.
@@
Figure 18: Slot selection for 4 slots. From left to right: (1) Linear action probes for VideoSaur slots pretrained basketball task for varying budget of trajectories. (2) Corresponding slot masks for basketball task after VideoSaur pretraining for number of slots $K=4$.
@@
Figure 19: Slot selection for 5 slots. From left to right: (1) Linear action probes for VideoSaur slots pretrained basketball task for varying budget of trajectories. (2) Corresponding slot masks for basketball task after VideoSaur pretraining for number of slots $K=5$.
@@
Figure 20: Slot selection for 8 slots. From left to right: (1) Linear action probes for VideoSaur slots pretrained basketball task for varying budget of trajectories. (2) Corresponding slot masks for basketball task after VideoSaur pretraining for number of slots $K=8$.
@@
Figure 21: Slot selection for 11 slots. From top to bottom: (1) Linear action probes for VideoSaur slots pretrained basketball task for varying budget of trajectories. (2) Corresponding slot masks for basketball task after VideoSaur pretraining for number of slots $K=11$.
@@
Figure 22: Slot selection for 15 slots. From top to bottom: (1) Linear action probes for VideoSaur slots pretrained basketball task for varying budget of trajectories. (2) Corresponding slot masks for basketball task after VideoSaur pretraining for number of slots $K=15$.
Appendix I Performance on aggregated slots
As shown in Figure 6, there is a noticeable drop in performance for $K=11$ and $K=15$ slots when selecting only the top-1 slot. Moreover, only the 11 slots and 15 slots configurations show non-zero performance on the top-2 slot. This suggests that meaningful information may be distributed across multiple slots rather than being concentrated in a single one. To investigate this further, we aggregated the two most relevant slots (the top-1 and top-2 slots) for both $K=11$ and $K=15$. The results of this aggregation experiment are presented in Figure 23. These results indicate that aggregating the top two slots indeed leads to better performance, compared to using them individually, supporting the hypothesis that useful information is spread across multiple slots.
Figure 23: Normalized success rates of the BC agent trained on latent actions for varying numbers of fine-tuning labeled trajectories. From left to right: performance with number of slots $K=11$; performance with number of slots $K=15$. TL;DR: Aggregation of top-2 slots for 11 and 15 slots, marked $\mathbf{11(I,II)}$ and, accordingly, $\mathbf{15(I,II)}$, gives higher performance compared to performance on individual slots, marked $\mathbf{11(I)}$, $\mathbf{11(II)}$, $\mathbf{11(III)}$ and $\mathbf{15(I)}$, $\mathbf{15(II)}$, $\mathbf{15(III)}$. Also we add a line $\mathbf{3(I)}$, which denotes the performance on the selected slot on 3 slots (same as in Figure 4), to both plots for reference.
@@ Appendix I Performance on aggregated slots
We begin with the STEVE model (Singh et al., 2022), a widely adopted and promising approach for object-centric learning. However, we found that STEVE struggled to accurately identify the main object in visual inputs. In Figure 25, we compare the slot projections of STEVE and VideoSAUR, highlighting key differences in their representations. We also note that STEVE is largely similar to SAVi (Kipf et al., 2022), differing primarily in its use of a transformer-based decoder instead of a pixel-mixture decoder.
In terms of performance, we evaluate the STEVE-based object-centric model on the hopper-hop task from DCS (Figure 24), which underscores the importance of a strong object-centric representation for achieving good final performance.
@@
Table 7: Hyperparameter optimization setup for latent action learning from images (used for lapo, lapo-masks).

### v3

@@ Appendix A Dataset collection
In this section, we provide further details on the data collection process for the two benchmark environments used in our experiments: the Distracting Control Suite (DCS) and the Distracting Meta-World (DMW). For each environment, we describe how the expert demonstration datasets were generated and outline their key properties.
Distracting Control Suite (DCS): The datasets were collected via expert policies trained on DCS via PPO (for cheetah-run, walker-run and hopper-hop) and SAC (for humanoid-walk). The scores of the experts are presented on the Table˜2. The final dataset of transitions for each task consists of 5k trajectories (1k transitions each). The observations in the dataset have a height and width of 64px. The same underlying expert trajectories were used for both the standard DCS (dynamic backgrounds only) and the DCS-Hard settings; the only difference is the visual distractions applied to the observations. The DCS-Hard setting combines all three available distractions: dynamic backgrounds, random color variations, and camera perturbations.
Distracting MetaWorld (DMW): The datasets were collected for the hammer, bin-picking, basketball, and soccer tasks using expert policies from MetaWorld. The scores of the experts are presented on the Table˜2. The final dataset of transitions for each task consists of 20k successful trajectories (we filtered out non-successful trajectories during dataset collection). The observations in the dataset have a height and width of 128px.
@@ Appendix B Training details
All experiments were run in H100 GPU 80GB. The models are trained in bfloat16 precision. Training duration is shown in Table˜4. We perform hyperparameter optimization for each model. The 3 best sets of parameters were evaluated and the best mean score across 3 seeds is reported in Figure˜4.
@@
Method DCS DMW
@@
Object-centric learning pretraining codebase was adopted from Videosaur Zadaianchuk et al. (2023). It utilizes DINOv2 (Oquab et al. 2023) pretrained encoder vit-base-patch14-dinov2.lvd142m from Timm (Wightman 2019) models hub. The images in the dataset are upscaled for dino encoder up to 518px. The hyperparameters for object-centric pretraining can be found in Table˜5
Latent action learning model for images and object-centric masks is formed from IDM and FDM models based on a ResNet-like CNN encoder and decoder. The hyperparameter optimization setups for latent action learning from images (used for lapo, lapo-masks) can be found in Table˜7.
Latent action learning model for object-centric slots is formed from IDM and FDM based on 3-layer MLP blocks with residual connections and GeLU activations to effectively process vector representations. The hyperparameter optimizations setups for latent action learning from representations (used for lapo-slots) can be found in Table˜7.
Appendix C Fixed Initialization for Slot Stability
$\texttt{Train:}\quad\boldsymbol{s}_{k}^{\text{(init)}}=\boldsymbol{\mu}_{k}+\boldsymbol{\sigma}_{k}\odot\boldsymbol{\epsilon},\quad\boldsymbol{\epsilon}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I}),\quad\texttt{Inference:}\quad\boldsymbol{s}_{k}^{\text{(inference)}}=\boldsymbol{\mu}_{k}.$
@@ Appendix D Evaluation scores on DCS-hard
We mark as DCS-hard, the experiments on Disctracting Control Suite with 3 types of distractions: dynamic backgroungs, color and camera position perturbations. The scale of color and camera variations is $0.1$. The aggregated mean scores for DCS-hard and the corresponding results on DCS (only dynamic backgrounds) are present in the Main Paper.
@@ Appendix E Correlation between probe and normalized success rates
We examine the correlation between the linear action probe and the evaluation performance, using the same budget of labeled trajectories for both probe training and supervised fine-tuning of object-centric latent action learning (LAPO-slots), on the basketball task from DMW.
As shown in Figure˜7, there is a strong correlation between the probe performance and the resulting success rate, suggesting that the probe can serve as a reliable method for slot selection.
Figure 7: Linear action probes and corresponding resulting normalized success rates of BC agents on same amounts of labeled trajectories for different number of slots (top-2 slots for $K\leq 5$, top-3 slots for $K\geq 8$). From top to bottom: (1) 32 labeled trajectories, (2) 64 labeled trajectories, (3) 128 labeled trajectories
@@ Appendix F Probe slots selection for DCS
Slot selection for tasks from the Distracting Control Suite: linear probes on different slots and corresponding visual masks examples. The slots with the lowest linear probe (top I) were used to obtain the results of tables and figures in the Main Paper.
@@ Appendix G Probe slots selection for DMW
Slot selection for tasks from the Distracting MetaWord: linear probes on different slots and corresponding visual masks examples. The slots with the lowest linear probe (top I) were used to obtain the results of tables and figures in the Main Paper.
@@ Appendix H Probe slots selection for different number of slots
Slot selection for varying number of slots for basketball task from Distracted MetaWord: linear probes on different slots, downstream performance and corresponding visual masks examples. The slots with the lowest linear probe (top I for Figures˜16, 17, 19 and 20 and top I+II for Figures˜21 and 22) were used to obtain the results Figure 5e in the Main Paper.
@@
Figure 16: Slot selection for 2 slots. From left to right: (1) Linear action probes for VideoSaur slots pretrained basketball task for varying budget of trajectories. (2) Downstream performance of top I and top II slots pretrained basketball task for varying budget of trajectories. (3) Corresponding slot masks for basketball task after VideoSaur pretraining for number of slots $K=2$.
@@
Figure 17: Slot selection for 3 slots. From left to right: (1) Linear action probes for VideoSaur slots pretrained basketball task for varying budget of trajectories. (2) Downstream performance of top I and top II slots pretrained basketball task for varying budget of trajectories. (3) Corresponding slot masks for basketball task after VideoSaur pretraining for number of slots $K=3$.
@@
Figure 18: Slot selection for 4 slots. From left to right: (1) Linear action probes for VideoSaur slots pretrained basketball task for varying budget of trajectories. (2) Downstream performance of top I and top II slots pretrained basketball task for varying budget of trajectories. (3) Corresponding slot masks for basketball task after VideoSaur pretraining for number of slots $K=4$.
@@
Figure 19: Slot selection for 5 slots. From left to right: (1) Linear action probes for VideoSaur slots pretrained basketball task for varying budget of trajectories. (2) Downstream performance of top I and top II slots pretrained basketball task for varying budget of trajectories. (3) Corresponding slot masks for basketball task after VideoSaur pretraining for number of slots $K=5$.
@@
Figure 20: Slot selection for 8 slots. From left to right: (1) Linear action probes for VideoSaur slots pretrained basketball task for varying budget of trajectories. (2) Downstream performance of top I and top II slots pretrained basketball task for varying budget of trajectories. (3) Corresponding slot masks for basketball task after VideoSaur pretraining for number of slots $K=8$.
@@
Figure 21: Slot selection for 11 slots. From left to right: (1) Linear action probes for VideoSaur slots pretrained basketball task for varying budget of trajectories. (2) Downstream performance of top I, top II, top III and concatenated top I+II slots pretrained basketball task for varying budget of trajectories. (3) Corresponding slot masks for basketball task after VideoSaur pretraining for number of slots $K=11$.
Figure 22: Slot selection for 15 slots. From left to right: (1) Linear action probes for VideoSaur slots pretrained basketball task for varying budget of trajectories. (2) Downstream performance of top I, top II, top III and concatenated top I+II slots pretrained basketball task for varying budget of trajectories. (3) Corresponding slot masks for basketball task after VideoSaur pretraining for number of slots $K=15$.
Appendix I Ablation on the usage of DINO features
Our primary method, LAPO-slots, leverages VideoSAUR, which utilizes a powerful DINOv2 pretrained encoder. This raises an important question: does the performance improvement stem from the object-centric slot representations, or simply from the stronger features provided by DINO? To disentangle these effects, we conduct an ablation study. We introduce a new baseline, LAPO-dino, which uses features of DINOv2’s CLS token from (Oquab et al. 2023). This allows for a direct comparison between applying DINO features to raw pixels (LAPO-dino) versus applying them within our object-centric framework (LAPO-slots). The results, shown in Figure˜23, demonstrate that DINO features alone do not yield a performance gain over standard LAPO, likely due to challenges specific to video processing.
Figure 23: Ablation study on the DMW basketball task to disentangle the effects of DINO features and object-centric slots. The plot compares four lines: LAPO (baseline), LAPO-clean (upper bound on clean data), LAPO-dino (LAPO based on DINO features), and LAPO-slots (our method, using DINO features within an object-centric framework).
@@ Appendix I Ablation on the usage of DINO features
We begin with the STEVE model (Singh et al. 2022), a widely adopted and promising approach for object-centric learning. However, we found that STEVE struggled to accurately identify the main object in visual inputs. In Figure˜25, we compare the slot projections of STEVE and VideoSAUR, highlighting key differences in their representations. We also note that STEVE is largely similar to SAVi (Kipf et al. 2022), differing primarily in its use of a transformer-based decoder instead of a pixel-mixture decoder. In terms of performance, we evaluate the STEVE-based object-centric model on the hopper-hop task from DCS (Figure˜24), which underscores the importance of a strong object-centric representation for achieving good final performance.
@@
Table 7: Hyperparameter optimization setup for latent action learning from images (used for LAPO, LAPO-masks, LAPO-clear).